NH4Cl + Ca(OH)2 --> NH3 + H2O + CaCl2

If 3.5 moles Ca(OH)2 are added to 380ml 1.3M NH4Cl how many moles of Ca(OH)2 will be left over?


1
Expert's answer
2021-01-29T08:19:47-0500
"2NH4Cl+Ca(OH)2\to2NH3+2H2O+CaCl"

1.3/2=n/1

n=1.3/2=0.65 mol


"\Delta n (Ca(OH)2)=3.5-0.65=2.85 mol"

So 2.85 mol Ca(OH)2 will be left over.